OpenStudy (anonymous):

ALGEBRA HELP PLEASE Simplify the product & Write in standard form. 20. -2n^2 (5n - 9 + 4n^2) =

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

What grade are you in?

OpenStudy (anonymous):

i cant do the standard form part but i can simplify/expand it?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

we use the FOIL method, and multiply it with each number, so -2n^2 x 5n = -10n^3 -2n^2 x -9 = 18n^2 -2n^2 x 4n^2 = -8n^4 put them all together, you get -10n^3+18n^2-8n^4

OpenStudy (anonymous):

Whats foil stand for?

OpenStudy (anonymous):

first, outer, inner, last

OpenStudy (anonymous):

@09chunga Thanks :)

OpenStudy (anonymous):

no prob ^^